OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

FIrst: Use (3, 7) and (−2, −8) to find the slope, m = (y2-y1)/(x2-x1)

OpenStudy (agent0smith):

Parallel lines have the same slope. Then, use point-slope form, y-y1 = m(x-x1) to find the equation of the line.

OpenStudy (anonymous):

wait please say that in a different way

OpenStudy (anonymous):

so I use the formula y = mx + b over here?

OpenStudy (agent0smith):

No... use (3, 7) and (−2, −8) to find the slope,\[\Large m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] That's the slope of the line, since parallel lines have the same slope. THEN once you have m, \[\Large y-y_1 = m(x-x_1)\] using the point (−2, 3) as (x1, y1)

OpenStudy (anonymous):

-15/-5? which is also 3

OpenStudy (agent0smith):

Good, that's m. Now work on the second part.

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Follow the steps, they're in my last post.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but which one will be y?

OpenStudy (agent0smith):

See the above post... you have m, and x1 and y1, now plug them into that equation. y will be y and x will be x.

OpenStudy (anonymous):

oh